OpenStudy (ny,ny):

I'm having trouble with slope and rate of change. Are they the same thing? How do you find them?

OpenStudy (amistre64):

they are related yes, not quite the same thing tho, but related

OpenStudy (amistre64):

slope is defined as:\[\frac{f(x+h)-f(x)}{(x+h)-h}\] this is also known as the average rate of change. there is a concept known as the instantaneous rate of change which is defined as:\[\lim_{h\to\ 0}\frac{f(x+h)-f(x)}{(x+h)-h}\]

OpenStudy (amistre64):

ugh, (x+h) - x is the understuff .... had a typo

OpenStudy (ny,ny):

so what's slope = rise/run

OpenStudy (amistre64):

thats a good way to look at it yes

OpenStudy (amistre64):

rise is the change in the value of y run is the change in the value of x

OpenStudy (amistre64):

if we rise by 2 as we run, or move, to the right by 7 slope is just 2/7

OpenStudy (amistre64):

if we drop by 3 as we move to the right by 4 the slope is defined as -3/4

OpenStudy (ny,ny):

so if i'm asked to find the slope of some graph thing... like  what do i do first. I have to do something with two points and then use \[slope=y ^{2}-y ^{1}/x ^{2}-x ^{1}\]... and yeah.... sorry i didn't understand in class. :P